OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4 y= 3x - 12 y= -3x y=-3x + 12 y= 3x

OpenStudy (anonymous):

Parallel lines have the same slope, do you know the slope?

OpenStudy (anonymous):

To be honest, I have NO clue.

OpenStudy (aroub):

y=mx+b where m is the slope here you have: y = -3x - 4 What do you think the slope ( m ) is?

OpenStudy (anonymous):

-3?

OpenStudy (aroub):

Bingo!

OpenStudy (anonymous):

Now you have the slope, were you able to find b?

OpenStudy (aroub):

you know the slope all you have to do now is find the y-intercept ( b ) : y=mx+b , (2, 6) , m=-3 ^ ^ x y Substitute; x by 2, y by 6 , m by -3 and you'll get b 6=-3(2) +b 6=-6 +b 12=b we got the y-intercept and the slope you can now form an equation :) y=-3x+12

OpenStudy (aroub):

How did you get 4 ?

OpenStudy (anonymous):

I have no idea..

myininaya (myininaya):

you could cheat on this since it is multiple choice you already eliminated two choices now the two choices you have not eliminated plug in (2,6) for both and see which equation holds

myininaya (myininaya):

or did what aroub did lol

myininaya (myininaya):

gj aroub

OpenStudy (anonymous):

Oh alright, well thank you everyone. It makes more sense now since it was explained like that.